OpenStudy (anonymous):

Fool's problem of the day, Just one easy Arithmetic problem, 1) Four friends went to a shop. Ramesh bought a gas stove and a glass and spent \( 720\$ \). Prakash bought a bucket and four glasses along with a towel and spend \(640\$\). Rahul and Mahesh spent a third of the total amount spent by four of them, and one of them bought only a towel and a bucket. How much was the gas stove worth, if Mahesh and Rahul spent an equal amount on their shopping? [Solved by @experimentX] Good luck!

OpenStudy (anonymous):

Hey! dont post your links in other's questions. its against the rules

OpenStudy (anonymous):

You can post the link in the chat below

OpenStudy (callisto):

Sorry, just want to make sure one thing, is total amount spent by them = amount spent by Pra. and Rame.?

OpenStudy (anonymous):

@experimentX: That is not the right answer.

OpenStudy (mani_jha):

Let the price of gas stove be s, glass be g, bucket be b and towel be t. s+g=720 b+4g+t=640 t+b=(1/2)*(720+640)/3 Put b+t in the second equation, get g and substitute that in the first equation. Then we can find s. Correct?

OpenStudy (anonymous):

whoa ,maini,good approach...

OpenStudy (callisto):

Ram : ST + G = 720 Pra : B +4G +T = 640 Rah + Ma: = 1/3 (720+640) where Rah = Ma and B+T = 1/2 [(1/3) (720+640)] B +T = 680/3 B +4G +T = 640 680/3 +4G = 640 G = (640 - 680/3)/4 = 310/3 ?

OpenStudy (anonymous):

yes

OpenStudy (mani_jha):

I am getting something close to 616.66. @sarkar, thanks.

OpenStudy (anonymous):

@Mani_Jha: What is your final answer?

OpenStudy (callisto):

Nah, Mani I don't mean to 'copy' you (well, i didn't )

OpenStudy (anonymous):

616.66

OpenStudy (anonymous):

thats Mani's final answer

OpenStudy (anonymous):

Okay, then it's not correct. The answer is an integer.

OpenStudy (mani_jha):

@Callisto, I know that you won't copy someone else's answer.

OpenStudy (anonymous):

s=(4040-x)/6 x=rahul or mahesh

OpenStudy (callisto):

is total amount spent by them = amount spent by Pra. and Rame.?

OpenStudy (anonymous):

I assumed all of them (4 people)..

OpenStudy (anonymous):

@Callisto: What do you think?

OpenStudy (callisto):

If I had known, I wouldn't have asked for twice

OpenStudy (mani_jha):

Total amount spent by Rahul and Mahesh= (1/3)*Amount spent by Prakash and Ramesh. That's what I think it is. Where could I've gone wrong!

OpenStudy (anonymous):

@Callisto: I have updated the problem statement. I hope it's clear now :)

OpenStudy (mani_jha):

Oh, that's where I went wrong!

OpenStudy (anonymous):

it is not good question, something is ambiguous

OpenStudy (turingtest):

A=s+g B=b+4g+t (C+D)=(A+B)/3 or (C+t+b=A+B)/3 3 equations, so it should be solvable

OpenStudy (anonymous):

Arithmetic is generally easy, the ambiguity is generally resolved by practice and experience.

OpenStudy (anonymous):

Yes it is. Congratz @experimentX

OpenStudy (turingtest):

whoa, and how did you get that @experimentX ?

OpenStudy (experimentx):

lol ... read the question clearly the latter guys spent 1/3 of the total .. means the former spent 2/3 ... which means we can find the total ... so it becomes easy peasy after that

OpenStudy (turingtest):

lol, yeah thanks I didn't see your post I just realized it while I was on another Q, got it now thx !